# Effective Backdoor Mitigation Depends on the Pre-training Objective

### v3

@@ 4 Analysis of Stronger Pre-training Objective
CleanCLIP successfully cleaned the models pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ by using $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$. However, it was unsuccessful for the model pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$. A plausible reason for this behavior could be that we are using the same pre-training and finetuning objective in the latter case, and CleanCLIP might be able to successfully clean these model if we choose to finetune with a distinct loss objective. To test this hypothesis, we finetune the models pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ using an additional deep clustering objective (deep-clustering-paper) on the vision encoder.
In deep clustering, we first obtain a pseudo-label for each image. A pseudo-label can be obtained either by performing a 1,000-way clustering on feature space of our trained vision encoder, using FAISS (faiss) or by using existing powerful models such as SigLIP ViT-L/14 (zero-shot Imagenet accuracy of 83.08%) siglip using zero-shot classification for each image into one of 1,000 Imagenet classes. In the case that we use clustering-based pseudo-label assignment, we repeat it every epoch. For every image in the cleaning dataset, we learn to predict the assigned pseudo-label ($\hat{y}$) with the help of a linear classifier on top of the vision encoder using cross-entropy loss ($\mathcal{L}_{\text{Xent}}$). Let $W\in\mathbb{R}^{d\times 1000}$ be the linear classifier mapping visual features ($\mathbb{R}^{d}$) to one of the 1000 pseudo-labels. For a given datapoint $(I_{j},T_{j})$ with assigned pseudo-label $\hat{y}_{j}$, deep clustering’s objective becomes
@@
and therefore overall objective becomes $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}+\mathcal{L}_{\text{DeepClust}}^{ft}$. In the case of clustering-based pseudo-label assignment, we re-initialize classifier head $W$ every time we re-compute pseudo-labels (similar to deep-clustering-paper). We emphasize that the use of SigLIP ViT-L/14 for obtaining pseudo-labels is a cheating experiment for a deep clustering task, however, this experiment is solely performed to probe the upper bound of the poison removal that can be obtained when using deep clustering approaches. We finetune the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model on the CC6M dataset for 10 epochs using 8 learning rates in both experiments. Similar to previous experiments, we measure the Top-1 zero-shot Imagenet accuracy and ASR at the end of each finetuning epoch.
Figure 4 shows the scatter plot of the two metrics when finetuning is performing using the additional deep clustering losses. As expected, we observe that the Pareto-frontier when using the pseudo-labels from SigLIP ViT-L/14 is better than the clustering-based pseudo-label assignment. However, neither of the two losses can successfully remove the poison. These experiments show that the suboptimality of CleanCLIP to clean the model pre-trained with a stronger objective is not just due to the same pre-training and finetuning loss, but due to stronger poison induction. Had the difference between the pre-training and finetuning objectives been the reason for CleanCLIP’s success, then our DeepClustering experiments could have removed the poison without sacrificing the clean accuracy. The experiments in the next sections further corroborate this observation.
Figure 4: Top-1 zero-shot Imagenet validation set accuracy v/s the ASR during the finetuning process for the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model on the CC6M dataset. The finetuning is done with $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}+\mathcal{L}^{ft}_{\text{DeepClust}}$. Takeaway: Adding $\mathcal{L}^{ft}_{\text{DeepClust}}$ is unable to successfully remove the poison from models pre-trained using the strong objective.
@@
heavy-regularization proposed heavy regularization as an approach to remove backdoors from vision and language models. To evaluate the efficacy of this approach, we finetune the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model on CC6M dataset with 9 regularization weight for 10 epochs using 8 learning rates for each weight. Figure 5 shows the scatter plot of the Top-1 zero-shot Imagenet validation set accuracy and the ASR at each epoch of the finetuning process. We observe that no hyperparameter combination can successfully remove the poison, which shows that simple regularization cannot remove the strongly induced poison.
Figure 5: Top-1 zero-shot Imagenet validation set accuracy v/s the ASR during the finetuning process for the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model on the CC6M dataset. The finetuning is done with each of the three aforementioned losses and varying regularization weights. Takeaway: Using regularization with any hyperparameter combinations is unable to remove the poison from models pre-trained using the strong objective.
@@
shrink-and-perturb suggested that warm starting models might hurt their finetuned performance. They recommend a perturbation approach in which a slight noise is added to the model weights. Specifically, $\theta_{i}^{t}\leftarrow\lambda\theta_{i}^{t-1}+p^{t}\text{, where }p^{t}\sim\mathcal{N}\left(0,\sigma^{2}\right)\text{ and }0<\lambda<1$. While this technique was developed to have better generalization with continued finetuning, we explore whether shrinking and perturbing can mitigate poisoning, as cleaning is done by finetuning. To evaluate whether this approach can help in removing the poison from the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model, we add noise with 5 noise scales ($\sigma$) for 15 values of $\lambda$. We measure the Top-1 zero-shot Imagenet validation set accuracy and the ASR of the obtained model and choose two models whose ASR was lower than 15% for finetuning. We finetune both the selected models using 8 learning rates for 20 epochs. Figure 6 shows the accuracy and ASR at the end of each finetuning epoch. We observe that even in this case none of the hyperparameter combinations can successfully remove the poison.
Figure 6: Top-1 zero-shot Imagenet validation set accuracy v/s the ASR during the finetuning process for the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model on CC6M dataset post shrinking and perturbing the weights. The finetuning is done with $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$. Takeaway: No hyperparameter combination is able to depoison the models pre-trained using the strong objective.
@@
We also perform ablation experiments to see the impact of hyperparameters on the ability of CleanCLIP to remove poison from the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model. In particular, we study the impact of the size of the cleaning dataset, the number of finetuning epochs, and the weightage of $\mathcal{L}^{ft}_{\text{SSL}}$. Appendix E reports the metrics when the finetuning is done on a dataset that is twice the size of the cleaning dataset used in the previous section. Appendix F reports the metrics when the finetuning is done for up to 100 epochs, i.e., 5$\times$ longer than in the previous section. Appendix G reports the accuracy and ASR metrics when the finetuning is done with higher weight to the $\mathcal{L}^{ft}_{\text{SSL}}$ loss. While changing the hyperparameters helps improve the Pareto-frontier curve, none of the three ablations can successfully remove the strongly induced poison.
@@
These experiments demonstrate that making the pre-training and finetuning objectives distinct from each other, and perturbation techniques like heavy regularization (heavy-regularization) and shrink-and-perturb (shrink-and-perturb) are not enough to remove the strongly induced poison. Moreover, we observe that while doing a more fine-grained search on different hyperparameters helps slightly improve the Pareto-frontier curve, none of the hyperparameter ablations can successfully remove the poison. Since removing the poison from such a model is an open research problem, a practitioner who is engaged in training models using web-curated data should consider training their model with only $\mathcal{L}^{pre}_{\text{MMCL}}$ so that they can clean their model on a carefully curated image-text pair dataset to remove potential backdoors from it, even when there are a few poisoned examples in the finetuning dataset.
5 Related Works
Contrastive Pre-training
Contrastive Learning was formally established in seminal works by bromley1993signature; chopra2005learning; hadsell2006dimensionality that has evolved, giving rise to contemporary algorithms such as CPC (oord2018representation), DCL (yeh2022decoupled), SimCLR (chen2020simple), and NNCLR (dwibedi2021little).  [^2] We refer the readers to balestriero2023cookbook for more development on SSL. These approaches, at their core, share a common objective: bringing similar elements (augmentation/retrieval) closer in representation space while pushing dissimilar ones apart.
radford2021learning extended this idea beyond a single modality to provide a dual-encoder approach for learning a shared representation space between image and text called CLIP. Images and their corresponding captions are brought close while the dissimilar images and captions are pushed away. jia2021scaling further extended this paradigm to handle noisy billion-scale datasets, demonstrating exceptional zero-shot accuracy across benchmarks like Imagenet-1K (5206848), MS-COCO retrieval, and robustness against variations in Imagenet-V2/R/A/C. Since then, there have been several improvements to the zero-shot accuracy, by adding components to the loss term. CyCLIP (goel2022cyclip) imposes additional consistency regularization; SLIP (mu2022slip) applies an additional self-supervision loss within image modality and was further unified by UniCLIP (lee2022uniclip). DeCLIP (li2021supervision) additionally uses kNN augmentation; FILIP (yao2021filip) additionally applies CLIP loss to fine-grained token representations. CLIP performance has also been improved by additional captioning loss (yu2022coca).
Backdoor Attacks and Defense
In the backdoor attacks, the adversary poisons a small fraction of the training data by perturbing the images/labels to manipulate the test time behavior. A prevalent form of this attack involves adding a trigger, such as a random pixel patch, into a small subset of the training dataset (souri2022sleeper; gu2017badnets; turner2018clean). During inference, models perform normally on images without the triggers but exhibit catastrophic failures when tested with the triggered images, erroneously predicting the labels targeted by the adversary. While the study of backdoor attacks has historically centered on supervised learning, recent attention has extended to self-supervised (saha2022backdoor) and multimodal representation learning (bansal2023cleanclip; carlini2021poisoning; carlini2023poisoning). This work focuses exclusively on the poisoning of multimodal models like the CLIP model.
The most common defense strategies against backdoor attacks primarily revolve around the identification and detection of poisoned examples (steinhardt2017certified; gao2019strip; wang2019neural; yang2022not; li2021antibackdoor; roclip; safeclip). However, alternative approaches have emerged, such as defense through knowledge distillation (yoshida2020disabling) and robust training procedures involving data augmentation (borgnia2021strong). Despite these efforts, research by carlini2021poisoning; carlini2023poisoning shows that poisoning even an exceedingly small fraction of the training datapoints (as little as 0.002%) can substantially impact model performance. Consequently, the effectiveness of detection-based methods in the context of multimodal pre-training remains uncertain. To address this challenge, bansal2023cleanclip propose CleanCLIP, a finetuning based procedure using a combination of MMCL and SSL losses, designed to cleanse the poisoned CLIP models, assuming access to a small, guaranteed to be a clean dataset.
Our Work
Our objective is to decipher the amenability of CleanCLIP to remove poison from pre-trained models under varying conditions like pre-training objective, lack of completely clean data, and lack of knowledge of the specific backdoor attack. Given that intramodal self-supervision loss has shown enhanced classification accuracy for multimodal models, we choose to investigate CleanCLIP effectiveness when models are trained with a combination of MMCL and SSL losses, vs. the case when they are just trained with MMCL loss. We also investigate it’s effectiveness when its finetuning data has a few poisoned datapoints and examine the stopping criterion for finetuning when the knowledge of the specific backdoor attack is not available.
@@ 6 Conclusions
Through our extensive hyperparameter search and ablation experiments, we unveil a critical limitation of the current state-of-the-art poison mitigation technique CleanCLIP. It fails to effectively remove backdoor poisoning when a model is pre-trained on stronger objectives like the combination of multimodal contrastive learning (MMCL) and intramodal self-supervised learning (SSL). This objective is common practice in popular approaches like SLIP (mu2022slip) and has demonstrated superior accuracy over pre-training with only the MMCL objective. Our experiments show that this vulnerability persists irrespective of the scale of the pre-training and the cleaning datasets.
Particularly concerning is the unstable cleaning trajectory in models using the stronger pre-training objective (Figure 1(b)). Practitioners, often unaware of the specific backdoor attack on their models, face challenges in determining the optimal point to halt the cleaning process. This instability can lead to suboptimal models, as continued finetuning can both decrease accuracy and increase attack success rate (ASR). Furthermore, our findings highlight the critical assumption of a completely poison-free cleaning dataset for CleanCLIP’s effectiveness, an assumption that is rarely met in practical scenarios. This becomes particularly problematic with the use of stronger pre-training objective. The models pre-trained with the simpler MMCL objective evade both these issues by having stable cleaning trajectories and amenabibility to poison reduction even under non-ideal conditions.
Given these insights, we urge practitioners to consider pre-training their models using the simpler MMCL objective. Even though this might slightly hurt the accuracy, it significantly enhances its amenability to remove backdoors. Our recommendation would also circumvent the issue of knowing when to halt the cleaning procedure, as more finetuning epochs would not hurt the model’s accuracy and ASR. Further, it will also be beneficial in scenarios where the cleaning data is not entirely poison-free.
Our work also underscores the formidable challenge of defending models against backdoor attacks, an open research problem. We invite the community to develop robust defense methods against stronger pre-training objectives.
Table 2: This table shows the original Top-1 zero-shot Imagenet validation set accuracy and ASR for the models pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ and $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ on the CC3M and CC6M datasets. These models are the input to CleanCLIP approach to remove their poison. Takeaway: The models pre-trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ achieve higher accuracy than their counterparts trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ solely. And adding poison to a mere 0.05% of the training data achieves almost a 100% ASR.
CC3M CC6M
Accuracy ($\uparrow$) ASR ($\downarrow$) Accuracy ($\uparrow$) ASR ($\downarrow$)
$\mathcal{L}^{pre}_{\text{MMCL}}$ 16.00% 99.88% 23.76% 99.98%
$\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ 17.04% 99.03% 23.86% 99.45%

### v4

@@ 5 Analysis of the Stronger Pre-training Objective
CleanCLIP successfully cleans the models trained with $\mathcal{L}^{pre}_{\text{MMCL}}$ by finetuning with $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$. However, it was unsuccessful for the model trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$. A plausible reason for this behavior could be that we are using the same pre-training and cleaning objective in the latter case, and CleanCLIP might be able to successfully clean the latter models if we were to clean it with a loss objective that is distinct from its pre-training objective. To test this hypothesis, we clean the models trained with $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ by finetuning with $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}+\mathcal{L}^{ft}_{\text{DeepClust}}$, where $\mathcal{L}^{ft}_{\text{DeepClust}}$ is an additional deep clustering objective (Caron et al., 2018) on the vision encoder.
In deep clustering, we first obtain a pseudo-label for each image. We obtain the pseudo-label for an image in two ways: a) by classifying each image into one of the 1,000 Imagenet classes using powerful models such as SigLIP ViT-L/14 (zero-shot Imagenet accuracy of 83.08%) (Zhai et al., 2023), and b) performing a 1,000-way clustering on feature space of our trained vision encoder, using FAISS (Johnson et al., 2019). Note that the use of SigLIP ViT-L/14 for obtaining pseudo-labels is a cheating experiment for a deep clustering task (we don’t have access to ground truth labels and therefore use SigLIP ViT-L/14 for this task); however, this experiment is solely performed to probe the upper bound of the poison removal that can be obtained when using deep clustering approaches.
Figure 7: Top-1 zero-shot Imagenet validation set accuracy vs. the ASR during the cleaning process for the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model on the CC6M dataset. The finetuning is done with $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}+\mathcal{L}^{ft}_{\text{DeepClust}}$. Takeaway: Adding $\mathcal{L}^{ft}_{\text{DeepClust}}$ is unable to successfully remove the poison from models pre-trained using the strong objective, indicating that having distinct pre-training and cleaning objectives does not ensure removal of poison.
In the latter case, when we use clustering-based pseudo-label assignment for every image in the cleaning dataset, we learn to predict the assigned pseudo-label ($\hat{y}$) with the help of a linear classifier on top of the vision encoder using cross-entropy loss ($\mathcal{L}_{\text{Xent}}$). Let $W\in\mathbb{R}^{d\times 1000}$ be the linear classifier mapping visual features ($\mathbb{R}^{d}$) to one of the 1000 pseudo-labels. For a given datapoint $(I_{j},T_{j})$ with assigned pseudo-label $\hat{y}_{j}$, deep clustering’s objective becomes
@@
and therefore overall objective becomes $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}+\mathcal{L}_{\text{DeepClust}}^{ft}$. Following Caron et al. (2018), we re-initialize classifier head $W$ every time we re-compute pseudo-labels. We finetune the model trained using $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ on the CC6M dataset for 10 epochs using 8 different learning rates for both the clustering techniques (see Appendix D for hyperparameter details) and measure the Top-1 zero-shot Imagenet accuracy and ASR at the end of each finetuning epoch.
Figure 7 shows the scatter plot of the Top-1 zero-shot Imagenet validation set accuracy and the ASR for this experiment. We observe that adding deep clustering does not successfully remove the poison from the model. This experiment shows that the ineffectiveness of CleanCLIP in cleaning the model trained with the stronger objective is not just due to the same pre-training and cleaning objectives but due to stronger poison induction. Had the difference between the pre-training and finetuning objectives been the reason for CleanCLIP’s success, then this DeepClustering experiment could have successfully cleaned the model.
@@
In this section, we attempt to remove the poison induced using $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ with heavy regularization. Zhu et al. (2023) proposed heavy regularization as an approach to remove backdoors from vision and language models. To evaluate the efficacy of this approach, we finetune these models with 9 different regularization weights for 10 epochs using 8 different learning rates for each regularization weight on 100K image-text pairs (see Appendix D for hyperparameter details). The regularization loss is added to three different loss objectives during the finetuning process: a) $\mathcal{L}^{ft}_{\text{MMCL}},b)\mathcal{L}^{ft}_{\text{SSL}}$, and c) $\mathcal{L}^{ft}_{\text{MMCL}}+\mathcal{L}^{ft}_{\text{SSL}}$.
Figure 9 shows the scatter plot of the Top-1 zero-shot Imagenet validation set accuracy and the ASR at the end of each finetuning epoch. We observe that no hyperparameter combination can successfully remove the poison, which shows that simply adding regularization cannot remove the strongly induced poison.
Figure 8: Top-1 zero-shot Imagenet validation set accuracy vs. the ASR during the cleaning process. Takeaway: Using regularization with any hyperparameter combinations is unable to remove the poison from models trained using $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$.
Figure 9: Top-1 zero-shot Imagenet validation set accuracy vs. the ASR during the cleaning process post shrinking and perturbing the weights. Takeaway: The shrink and perturb technique is unable to remove poison from the model that is trained using $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$.
@@
In this section, we attempt to remove the poison using the shrink and perturb technique. Ash & Adams (2020) proposed shrink and perturb technique to improve the accuracy of a model when finetuned for a task. CleanCLIP also cleans a model by finetuning, and therefore, it could benefit from this technique. In this technique, a small noise is added to the model weights before starting finetuning, $\theta_{0}\leftarrow\lambda\theta_{0}+p_{0}\text{, where }p\sim\mathcal{N}\left(0,\sigma^{2}\right)\text{ and }\lambda\in(0,1)$. To evaluate whether this approach can help in removing the poison from a model trained using $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$, we add a small noise to all its weights. While the choice of noise scale ($\sigma$) and shrinking parameter ($\lambda$) is a hyperparameter, we experiment with 5 values of $\sigma$ for 15 values of $\lambda$. To determine which model to clean amongst the 75 possible models obtained from shrink and perturb described above, we measure the Top-1 zero-shot Imagenet validation set accuracy and the ASR of the models with noised weights, and selected model with the highest accuracy whose ASR was lower than 15% for cleaning. We clean the selected model by finetuning it using 8 learning rates for 20 epochs on 100K image-text pairs. Figure 9 shows the accuracy and ASR at the end of each finetuning epoch. We observe that even this approach is unable to successfully remove the poison from the models.
@@
We also perform ablation experiments to see the impact of hyperparameters on the ability of CleanCLIP to remove poison from the $\mathcal{L}^{pre}_{\text{MMCL}}+\mathcal{L}^{pre}_{\text{SSL}}$ pre-trained model. In particular, we study the impact of the size of the cleaning dataset, the number of finetuning epochs, and the weightage of $\mathcal{L}^{ft}_{\text{SSL}}$. Appendix G reports the metrics when the finetuning is done on a dataset that is twice the size of the cleaning dataset used in the previous section. Appendix H reports the metrics when the finetuning is done for up to 100 epochs, i.e., 5$\times$ longer than in the previous section. Appendix I reports the accuracy and ASR metrics when the finetuning is done with higher weight to the $\mathcal{L}^{ft}_{\text{SSL}}$ loss. While changing the hyperparameters helps improve the Pareto-frontier curve, none of the three ablations can successfully remove the strongly induced poison.
@@
These experiments demonstrate that making the pre-training and finetuning objectives distinct from each other and perturbation techniques like heavy regularization (Zhu et al., 2023) and shrink-and-perturb (Ash & Adams, 2020) are not enough to remove the strongly induced poison. Moreover, we observe that while doing a more fine-grained search on different hyperparameters helps slightly improve the Pareto-frontier curve, none of the hyperparameter ablations can successfully remove the poison. Since removing the poison from such a model is an open research problem, a practitioner who is engaged in training models using web-curated data should consider training their model with only $\mathcal{L}^{pre}_{\text{MMCL}}$ so that they can clean their model on a carefully curated image-text pair dataset to remove potential backdoors from it, even when there are a few poisoned examples in the finetuning dataset.
@@ 6 Conclusions
Through our extensive hyperparameter search and ablation experiments, we unveil a critical limitation of the current state-of-the-art poison mitigation technique for multimodal models, CleanCLIP. It fails to effectively remove backdoor poisoning when a model is trained using stronger objectives like the combination of multimodal contrastive learning (MMCL) and intramodal self-supervised learning (SSL). This objective is common in popular approaches like SLIP (Mu et al., 2022) and has demonstrated superior accuracy over training with only the MMCL objective. Our experiments show that this vulnerability persists irrespective of the scale of the pre-training and the cleaning datasets, irrespective of the manner of poison induction (from scratch or by finetuning), and irrespective of the specific backdoor attack.
Particularly concerning is the unstable cleaning trajectory in models trained using the stronger objective (Figure 5(b)). Often unaware of the specific backdoor attack, practitioners face challenges in determining the optimal point to halt the cleaning process. This instability can lead to suboptimal models, as continued finetuning can decrease accuracy and increase attack success rate (ASR). Furthermore, our findings highlight the critical assumption of a completely poison-free cleaning dataset for CleanCLIP’s effectiveness, an assumption that is rarely met in practical scenarios. This becomes particularly problematic with the use of stronger pre-training objectives. The models trained with the simpler MMCL objective evade both these issues by having stable cleaning trajectories and amenability to poison reduction even under non-ideal conditions.
Given these insights, we urge practitioners to consider training their models using the simpler MMCL objective. Even though this might slightly hurt the accuracy, it significantly enhances its amenability to remove backdoors. Our recommendation would also circumvent the issue of knowing when to halt the cleaning procedure, as more finetuning epochs would not hurt the model’s accuracy and ASR. Further, it will also be beneficial when cleaning data is not entirely poison-free. Our work underscores the formidable challenge of defending models against backdoor attacks, an open research problem. We encourage future defense methods to robustly test their approach against various pre-training objectives and we invite the community to develop robust defense methods.